Find the gradient of a scalar function of position V where V (x, y, z) = x2y2 + ez, Calculate the magnitude of gradient at point P(1,5,-2)


1
Expert's answer
2021-02-07T19:07:57-0500

Answer

Gradient of V

"\nabla V=(i\frac{dV}{dx}+j\frac{dV}{dy}+k\frac{dV}{dz})"

"=i2xy^2+j2x^2y+ke^z"

At point P(1,5,-2)

"\nabla V= 50i+10j+ke^{-2}"